# ViewSpatial-Bench: Evaluating Multi-perspective Spatial Localization in Vision-Language Models

### v1

@@ 2 Related Works
Recently, VLMs have demonstrated significant advancements in understanding and reasoning about visual content [13, 14, 15]. Both proprietary and open-source models have achieved impressive performance in visual question answering (VQA), image captioning, and complex multimodal reasoning tasks. These models typically employ Transformer [16] architectures, incorporating image encoders and vision-language fusion modules [17, 18, 19], and are pre-trained on large-scale image-text pairs to establish strong connections between visual elements and linguistic descriptions [20].
However, despite current VLMs’ exceptional performance on certain visual reasoning tasks, their spatial understanding capabilities remain fundamentally limited [2, 21]. When handling tasks involving spatial relationships, object localization, or embodied interaction reasoning, models typically rely on camera-centric reference frames, with their spatial understanding strictly bound to the observational viewpoint [4, 22]. This constraint impairs their generalization capabilities and practical utility in tasks requiring perspective transformation or third-person spatial comprehension. Consequently, developing models with stronger perspective-taking awareness has emerged as a critical challenge for advancing multimodal intelligence to higher levels of sophistication.
@@
Although some benchmarks have begun to address cross-viewpoint spatial understanding, such as 3DSRBench [25] and SPHERE [26], which evaluate existing VLMs across various spatial tasks, they remain insufficient in terms of multi-task comprehensiveness and depth of perspective transformation assessment.
@@ 3.1 Overview
ViewSpatial-Bench comprises five localization recognition tasks across two complementary perspective frameworks. From the camera perspective: (1) Object Relative Direction recognition(Cam-Rel. Dir.), which determines spatial relationships between objects directly from images; (2) Object View Orientation recognition(Cam-Obj. Oir.), which identifies the gaze direction of individuals relative to the camera from an egocentric viewpoint. These tasks evaluate VLMs’ intuitive, egocentric spatial understanding abilities. From the human perspective: (3) Object Relative Direction recognition(Per-Rel. Dir.), which involves adopting the viewpoint of a character in the image to determine the spatial relationships of other objects from their perspective; (4) Object View Orientation recognition(Per-Obj. Oir.), which requires assuming the position of a character in the image to determine the direction of their gaze; (5) Scene Simulation Relative Direction recognition(Per-Sce. Sim.), which requires modeling oneself within a spatial scene across sequential frames to determine relative positions of other objects. These latter three tasks assess VLMs’ abstract, perception-dependent spatial awareness while accommodating complex human pose variations and spatial information in embodied scenarios.
Figure 2: ViewSpatial-Bench construction pipeline. From data collection to QA generation across camera perspective ( ) and human perspective ( ) tasks. The pipeline includes metadata creation, automatic filtering, spatial relation extraction, and manual verification.
@@ 3.2 Dataset Construction
ViewSpatial-Bench construction follows a systematic process using two complementary data sources: ScanNet for rich 3D scene reconstructions with accurate spatial coordinates, and MS-CoCo for diverse images with human subjects and annotated keypoints. This combination supports both precise 3D spatial reasoning and perspective-dependent human-centric understanding tasks. We developed specialized processing pipelines for each source to extract reliable spatial relationships using automated techniques with manual verification.
@@
For Cam-Obj. Oir. and Per-Obj. Oir. tasks, plus Per-Rel. Dir. task, we utilized the MS-CoCo validation set. We filtered images containing animate objects occupying at least 20% of the image area.
@@
0:  Image $I$, keypoints $K$, bounding box $B$, Orient-Anything model $D$
@@
To ensure the quality of ViewSpatial-Bench, we implemented a multi-stage filtering process for all tasks. During metadata generation, we eliminated invalid data with incorrectly calculated orientation angles or excessively large rotation angles. In the manual filtering stage, for relative direction tasks, we removed instances where objects were too close to each other, objects were difficult to identify, or images were blurry. For gaze direction recognition tasks, we filtered out data where subjects’ gaze directions significantly differed from their head orientations or where subjects were difficult to identify. Following automated construction and filtering, we conducted manual verification to confirm that target objects were clearly visible in images and that the spatial localizations were correct and unambiguous. This iterative refinement process continued until ViewSpatial-Bench met our quality standards [22, 23]. More dataset details are in Appendix B.
@@ 3.3 Dataset Statistics
Figure 4 shows the frequency distribution of spatial prepositions and objects in ViewSpatial-Bench. As illustrated in Figure 4(a), our benchmark incorporates a comprehensive set of directional terms, with balanced representation of primary directions ("front", "back", "left", "right") and compound directions ("front-left", "back-right", etc.). This diverse coverage ensures thorough evaluation of VLMs’ ability to process complex spatial relationships from multiple perspectives. Figure 4(b) depicts the distribution of the top 20 objects in ViewSpatial-Bench. The object distribution reflects common entities encountered in everyday environments, with furniture items (chair, table, sofa, desk) and personal objects (person, cup, bottle) well represented. This ensures practical relevance of the benchmark to real-world spatial reasoning scenarios, particularly for embodied AI applications that must navigate and interact with common objects.
(a) Answer Direction Distribution
(b) Top 20 Objects Frequency
@@ 4 Multi-View Spatial Model
To address the limitations in perspective-dependent spatial reasoning identified in current VLMs, we developed the Multi-View Spatial Model (MVSM) through a systematic enhancement approach. Our methodology combines high-quality training data with a specialized fine-tuning strategy designed specifically for multi-viewpoint spatial understanding. Following the ViewSpatial-Bench construction pipeline, we leveraged our automated spatial annotation framework to generate approximately 43K diverse spatial relationship samples across all five task categories. This dataset incorporates 3D spatial information from ScanNet [27] and MS-COCO [28] training sets, supplemented with Spatial-MM [4] data for the Per-Rel. Dir. task where full automation proved challenging due to complex human spatial coordinates and environmental contexts. Our Multi-Perspective Fine-Tuning strategy explicitly trains the model to reason from different observational viewpoints, enabling MVSM to develop a more unified representation of 3D spatial relationships that supports robust reasoning across both camera and human perspectives.
@@ 5.1 Experimental Setup
For our fine-tuning experiments, we use Qwen2.5-VL [37] as the backbone model. Following standard practice in efficient adaptation, we freeze the vision encoder and multi-modal projector while keeping the language model trainable. The model is trained for 3 epoch with an effective batch size of 16, achieved through gradient accumulation (4 steps with per-device batch size of 1) across 4 NVIDIA A100 (40GB) GPUs. The entire training process requires approximately 8.5 GPU hours, making our approach computationally efficient and accessible.More model training and evaluation details are shown in Appendix C.
@@ 5.3.1 Transfer Learning Performance
VSI-Bench VSI-App Model Rel Dir Route Plan Average Indoor Outdoor Average Qwen2.5-VL(3B)[37] 46.00 21.90 41.97 24.00 20.00 22.00 MVSM 46.93 $\uparrow$ 0.93 31.44 $\uparrow$ 9.54 44.34 $\uparrow$ 2.37 44.00 $\uparrow$ 20.00 24.00 $\uparrow$ 4.00 34.00 $\uparrow$ 12.00
@@ Appendix A Limitations
Our Camera-perspective Relative Direction tasks utilize exclusively indoor environments from ScanNet, potentially limiting generalizability to outdoor settings. As our transfer learning experiments on VSI-App suggest, there exists a substantial domain gap between indoor and outdoor spatial reasoning tasks. Outdoor environments present different spatial scales, object densities, and visual characteristics that may require specialized training approaches beyond those presented in this work.
@@
ViewSpatial-Bench evaluates only static spatial orientation comprehension without addressing dynamic spatial reasoning scenarios where objects or observers move through environments. Such dynamic reasoning represents an important aspect of embodied spatial cognition relevant to many practical applications, including robot navigation and interactive systems. Extending our benchmark to incorporate temporal sequences and motion-based spatial reasoning would provide a more comprehensive evaluation framework for embodied AI systems.
@@ B.1 Dataset Collection and Unification
Based on MS-COCO dataset annotations, we filter samples containing biological objects that occupy at least 20% of the image area to ensure sufficient visual salience of target objects. We subsequently employ manual annotation to filter out samples where gaze direction significantly deviates from head orientation, ensuring consistency in spatial orientation labeling. The filtered samples are then processed by the Orient-Anything-Large model for automatic head and body orientation angle annotation. Given that this model exhibits labeling errors when processing low-resolution images or objects with ambiguous directional tendencies, we conduct focused manual verification and data correction on extreme angle samples (excessively large or small angles). This quality assurance mechanism ensures the annotation accuracy of the final dataset.
@@ B.2 Data Statiscs
As shown in the word cloud analysis in Figure 6, our dataset is primarily constructed around two major categories: humans and objects, which aligns with our dual spatial localization task design targeting both camera and human perspectives. Tabl 4 provides a detailed breakdown of sample distributions across different task types in ViewSpatial-Bench.
@@ B.4 VSI-App Dataset Construction
For the ViewSpatial Interaction Application Dataset (VSI-App), we employ a three-stage human curation approach to construct a dataset specifically designed to evaluate multi-view spatial models (MVSM) capabilities in spatial reasoning for human-computer interaction under Out-of-Distribution scenarios. Initially, two professional annotators carefully screened and downloaded 50 high-quality scene images from professional online image platforms, with 25 indoor and 25 outdoor scenes respectively. Image selection strictly adheres to the following criteria: scenes must be highly consistent with indoor/outdoor themes, contain rich three-dimensional spatial hierarchical information, include clearly identifiable human subjects as viewpoint references, and demonstrate explicit spatial relationships and potential interaction possibilities between humans and other objects in the scene. This meticulous scene selection ensures that the dataset can adequately simulate the complex spatial environments of real-world human-computer interactions.
@@ C.1 Implementation Details
We select Qwen2.5-VL as the base model for supervised fine-tuning. The Cam-Rel. Dir., Cam-Obj. Dir., Per-Obj. Dir., and Per-Sce. Sim. tasks in the training dataset are generated through our automated construction pipeline using unified QA templates. The Per-Rel. Dir. task is constructed based on the Spatial-MM dataset, with language models employed to polish questions and enhance sample diversity. The distribution of training samples across tasks is detailed in Table 4.
@@ C.2 Evaluation Details
We evaluate all models under zero-shot settings, where models must directly predict the correct option based on given images and questions. Accuracy is calculated by comparing model predictions with ground truth answers. The prompt template used for evaluation is shown below.
@@
We follow the original paper’s experimental settings for VSI-Bench evaluation. We employ the lmms-eval framework to conduct zero-shot testing with a batch size of 1 and maximum frame count set to 32. All models are evaluated on a single GPU environment (A6000 48G) using the accelerate launcher.
@@
Figure 7: ViewSpatial-Bench Examples (Part1). Performance comparison of three models (Qwen2.5-VL(3B), GPT-4o, and MVSM) on five spatial reasoning tasks from camera perspective ( ) and human perspective ( ).

### v2

# Dingming Li1,2,*Hongxing Li1,*Zixuan Wang1Yuchen Yan1 Hang Zhang1Siqi Chen1Guiyang Hou1Shengpei Jiang3

@@ 2 Related Works
Recently, VLMs have demonstrated significant advancements in understanding and reasoning about visual content [13, 14]. Both proprietary and open-source models have achieved impressive performance in visual question answering, image captioning, and complex multimodal reasoning tasks. These models typically incorporate image encoders and vision-language fusion modules [15, 16, 17], pre-trained on large-scale image-text pairs [18].
However, despite current VLMs’ exceptional performance on certain visual reasoning tasks, their spatial understanding capabilities remain fundamentally limited [2, 4]. When handling tasks involving spatial relationships, object localization, or embodied interaction reasoning, models typically rely on camera-centric reference frames, with their spatial understanding strictly bound to the observational viewpoint [4, 19]. This constraint impairs their generalization capabilities and practical utility in tasks requiring perspective transformation or third-person spatial comprehension, making the development of models with stronger perspective-taking awareness a critical challenge for advancing multimodal intelligence.
@@
Although some benchmarks have begun to address cross-viewpoint spatial understanding, such as 3DSRBench [22] and SPHERE [23], they remain insufficient in terms of multi-task comprehensiveness and depth of perspective transformation assessment.
@@ 3.1 Overview
ViewSpatial-Bench comprises five localization recognition tasks across two complementary perspective frameworks. From the camera perspective: (1) Object Relative Direction recognition(Cam-Rel. Dir.), which determines spatial relationships between objects directly from images; (2) Object View Orientation recognition(Cam-Obj. Ori.), which identifies the gaze direction of individuals relative to the camera from an egocentric viewpoint. These tasks evaluate VLMs’ intuitive, egocentric spatial understanding abilities. From the human perspective: (3) Object Relative Direction recognition(Per-Rel. Dir.), which involves adopting the viewpoint of a character in the image to determine the spatial relationships of other objects from their perspective; (4) Object View Orientation recognition(Per-Obj. Ori.), which requires assuming the position of a character in the image to determine the direction of their gaze; (5) Scene Simulation Relative Direction recognition(Per-Sce. Sim.), which requires modeling oneself within a spatial scene across sequential frames to determine relative positions of other objects. These latter three tasks assess VLMs’ abstract, perception-dependent spatial awareness while accommodating complex human pose variations and spatial information in embodied scenarios.
Figure 2: ViewSpatial-Bench construction pipeline. From data collection to QA generation across camera perspective ( ) and person perspective ( ) tasks. The pipeline includes metadata creation, automatic filtering, spatial relation extraction, and manual verification.
@@ 3.2 Dataset Construction
ViewSpatial-Bench construction follows a systematic process using two complementary data sources: ScanNet for rich 3D scene reconstructions with accurate spatial coordinates, and MS-CoCo for diverse images with human subjects and annotated keypoints. This combination supports both precise 3D spatial reasoning and perspective-dependent person-centric understanding tasks. We developed specialized processing pipelines for each source to extract reliable spatial relationships using automated techniques with manual verification.
@@
For Cam-Obj. Ori. and Per-Obj. Ori. tasks, plus Per-Rel. Dir. task, we utilized the MS-CoCo validation set. We filtered images containing animate objects occupying at least 20% of the image area.
@@
0:  Image II, keypoints $K$, bounding box $B$, Orient-Anything model $D$
@@
To ensure the quality of ViewSpatial-Bench, we implemented a multi-stage filtering process for all tasks. During metadata generation, we eliminated invalid data with incorrectly calculated orientation angles or excessively large rotation angles. In the manual filtering stage, for relative direction tasks, we removed instances where objects were too close to each other, objects were difficult to identify, or images were blurry. For gaze direction recognition tasks, we filtered out data where subjects’ gaze directions significantly differed from their head orientations or where subjects were difficult to identify. Following automated construction and filtering, we conducted manual verification to confirm that target objects were clearly visible in images and that the spatial localizations were correct and unambiguous. This iterative refinement process continued until ViewSpatial-Bench met our quality standards [19, 20]. Additional dataset construction details are provided in Appendix B.1.
@@ 3.3 Dataset Statistics
Table 1 presents a comprehensive comparison between ViewSpatial-Bench and existing spatial reasoning benchmarks. ViewSpatial-Bench contains 5,712 samples across 1,338 unique scenes, employing a hybrid construction method that combines automated 3D annotation pipelines with manual verification. The benchmark distinguishes itself with 18 distinct directional categories and precise 3D coordinate annotations from ScanNet, while uniquely supporting evaluation across both camera-perspective and human-perspective viewpoints for comprehensive assessment of perspective-taking capabilities essential for embodied AI applications. Detailed statistical analysis is provided in Appendix B.2.
Benchmark
Construct
Method
Visual
Diversity
Scale & Diversity 3D Annotation Multi-Perspective Spatial Query Target
Samples Scenes Directions 3D-Coord Camera Person Person-Target Object-Target
SpatialRGPT-Bench [2] Automated Single 1,410 524 6
EmbSpatial-Bench [20] Automated Single 3,640 284 6
What’sUP [21] Manual Single 820 205 12
VSI-Bench [19] Automated Multi 3,672 245 8
3DSRBench [22] Manual Single 2,772 1,827 8
SPHERE [23] Manual Single 2,285 1,001 7
All-Angles Bench [28] Hybrid Single 2,132 90 4
GSR-BENCH [29] Automated Multi 820 205 12
ViewSpatial-Bench Hybrid Multi 5,712 1,338 18
Table 1: Comparison of ViewSpatial-Bench with existing spatial reasoning benchmarks. ViewSpatial-Bench provides the first comprehensive evaluation framework for multi-perspective spatial localization, uniquely supporting both camera and person viewpoints with the broadest scope of directional categories and spatial query targets.
@@ 4 Multi-View Spatial Model
To address the limitations in perspective-dependent spatial reasoning identified in current VLMs, we developed the Multi-View Spatial Model (MVSM) through a systematic enhancement approach. Our methodology combines high-quality training data with a specialized fine-tuning strategy designed specifically for multi-viewpoint spatial understanding. Following the ViewSpatial-Bench construction pipeline, we leveraged our automated spatial annotation framework to generate approximately 43K diverse spatial relationship samples across all five task categories. This dataset incorporates 3D spatial information from ScanNet [24] and MS-COCO [25] training sets, supplemented with Spatial-MM [4] data for the Person-perspective Relative Direction task where full automation proved challenging due to complex human spatial coordinates and environmental contexts. We structured all training data as image-text pairs using consistent natural language templates to articulate spatial relationships between objects or entities, with answers represented as standardized directional classifications. Our Multi-Perspective Fine-Tuning strategy explicitly trains the model to reason from different observational viewpoints, enabling MVSM to develop a more unified representation of 3D spatial relationships that supports robust reasoning across both camera and human perspectives.
@@ 5.1 Experimental Setup
We employ our MVSM fine-tuning strategy on Qwen2.5-VL-3B [36] as the backbone model. Detailed training configurations and evaluation procedures are provided in Appendix C.1 and  C.2.
@@ 5.3.1 Transfer Learning Performance
VSI-Bench VSI-App Model Rel Dir Route Plan Average Indoor Outdoor Average GPT-4o [37] 41.30 31.50 39.66 34.00 27.00 30.50 Qwen2.5-VL(3B) [36] 46.00 21.90 41.97 18.00 27.00 22.50 MVSM 46.93 $\uparrow$ 0.93 31.44 $\uparrow$ 9.54 44.34 $\uparrow$ 2.37 41.00 $\uparrow$ 23.00 36.00 $\uparrow$ 9.00 38.50 $\uparrow$ 16.00
@@ Appendix A Limitations
Our Camera-perspective Relative Direction tasks utilize exclusively indoor environments from ScanNet [24], potentially limiting generalizability to outdoor settings. As our transfer learning experiments on VSI-App suggest, there exists a substantial domain gap between indoor and outdoor spatial reasoning tasks. Outdoor environments present different spatial scales, object densities, and visual characteristics that may require specialized training approaches beyond those presented in this work.
@@
ViewSpatial-Bench evaluates only static spatial orientation comprehension without addressing dynamic spatial reasoning scenarios where objects or observers move through environments. Such dynamic reasoning represents an important aspect of embodied spatial cognition relevant to many practical applications, including robot navigation and interactive systems [39]. Extending our benchmark to incorporate temporal sequences and motion-based spatial reasoning would provide a more comprehensive evaluation framework for embodied AI systems.
@@ B.1 Dataset Collection and Unification
Based on MS-CoCo [25] dataset annotations, we filter samples containing biological objects that occupy at least 20% of the image area to ensure sufficient visual salience of target objects. We subsequently employ manual annotation to filter out samples where gaze direction significantly deviates from head orientation, ensuring consistency in spatial orientation labeling. The filtered samples are then processed by the Orient-Anything-Large model for automatic head and body orientation angle annotation. Given that this model exhibits labeling errors when processing low-resolution images or objects with ambiguous directional tendencies, we conduct focused manual verification and data correction on extreme angle samples (excessively large or small angles). This quality assurance mechanism ensures the annotation accuracy of the final dataset.
@@ B.2 Data Statiscs
As shown in the word cloud analysis in Figure 5, our dataset is primarily constructed around two major categories: humans and objects, which aligns with our dual spatial localization task design targeting both camera and human perspectives. Table 5 provides a detailed breakdown of sample distributions across different task types in ViewSpatial-Bench.
Figure 6 shows the frequency distribution of spatial prepositions and objects in ViewSpatial-Bench. As illustrated in Figure 4(a), our benchmark incorporates a comprehensive set of directional terms, with primary directions ("front", "right", "left") showing higher frequency than compound directions ("front-left", "back-right", "above-left"). This diverse coverage ensures thorough evaluation of VLMs’ ability to process complex spatial relationships from multiple perspectives, reflecting the natural usage patterns of spatial language.
Figure 6(b) depicts the distribution of the top 20 objects in ViewSpatial-Bench. The object distribution reflects common entities encountered in everyday environments, with furniture items (chair, table, sofa, desk) and personal objects well represented. This ensures practical relevance of the benchmark to real-world spatial reasoning scenarios, particularly for embodied AI applications that must navigate and interact with common objects.
(a) Answer Direction Distribution
(b) Top 20 Objects Frequency
@@ B.4 VSI-App Dataset Construction
For the ViewSpatial Interaction Application Dataset (VSI-App), we employ a three-stage human curation approach to construct a dataset specifically designed to evaluate multi-view spatial models (MVSM) capabilities in spatial reasoning for human-computer interaction under Out-of-Distribution scenarios. Initially, two professional annotators carefully screened and downloaded 200 high-quality scene images from professional online image platforms, with 100 indoor and 100 outdoor scenes respectively. Image selection strictly adheres to the following criteria: scenes must be highly consistent with indoor/outdoor themes, contain rich three-dimensional spatial hierarchical information, include clearly identifiable human subjects as viewpoint references, and demonstrate explicit spatial relationships and potential interaction possibilities between humans and other objects in the scene. This meticulous scene selection ensures that the dataset can adequately simulate the complex spatial environments of real-world human-computer interactions.
@@ C.1 Implementation Details
We select Qwen2.5-VL-3B [36] as the base model for supervised fine-tuning. The Cam-Rel. Dir., Cam-Obj. Ori., Per-Obj. Ori., and Per-Sce. Sim. tasks in the training dataset are generated through our automated construction pipeline using unified QA templates. The Per-Rel. Dir. task is constructed based on the Spatial-MM [4] dataset, with language models employed to polish questions and enhance sample diversity. The distribution of training samples across tasks is detailed in Table 5.
Following standard practice in efficient adaptation, we freeze the vision encoder and multi-modal projector while keeping the language model trainable. The model is trained for 3 epoch with an effective batch size of 16, achieved through gradient accumulation (4 steps with per-device batch size of 1) across 4 NVIDIA A100 (40GB) GPUs. The entire training process requires approximately 8.5 GPU hours, making our approach computationally efficient and accessible.
@@ C.2 Evaluation Details
We evaluate all models under zero-shot settings, where models must directly predict the correct option based on given images and questions. For API-based models, we used their standard online interfaces with default parameters. For open-source models, we employed their default generation settings through the Transformers [40] library. To ensure consistency and reliability in our results, each model was evaluated five times. The results reported in the manuscript represent the average performance across these multiple runs. Accuracy is calculated by comparing model predictions with ground truth answers. The prompt template used for evaluation is shown below.
@@
We follow the original paper’s experimental settings for VSI-Bench [19] evaluation. We employ the lmms-eval framework to conduct zero-shot testing with a batch size of 1 and maximum frame count set to 32. All models are evaluated on a single GPU environment (A6000 48G) using the accelerate launcher.
@@
C.3 Analysis experiment
Experiment Backbone Model Original MVSM Improvement
Training Format Qwen2.5-VL (3B) 35.85% 82.09% (MC) +46.24%
79.34% (DA) +43.49%
Multi-Backbone Qwen2.5-VL (3B) 35.85% 82.09% +46.24%
Qwen2.5-VL (7B) 36.85% 83.01% +46.16%
InternVL2.5 (2B) 34.98% 76.45% +41.47%
Table 6: Comprehensive analysis of MVSM training robustness across different question formats and model architectures. MC denotes Multiple Choice format, DA denotes Direct Answer format.
Training Format and Shortcut Learning Analysis. To verify that our Multi-View Spatial Model’s performance improvements stem from genuine spatial reasoning rather than shortcut learning, we conducted controlled experiments comparing different training formats. We used the same multi-perspective spatial dataset described in Section 4, containing approximately 43K samples across all five task categories. To eliminate potential shortcut learning through option elimination strategies, we converted the original multiple-choice format into a direct answer format where models generate spatial directions without candidate options.
Following identical experimental settings as described in Section 5.1, we trained Qwen2.5-VL(3B) on both formats. As shown in table 6, both approaches yielded substantial improvements over the baseline, with multiple-choice format achieving 82.09% overall accuracy compared to 79.34% for direct answer format. The minimal performance difference between formats confirms that MVSM’s gains result from enhanced spatial understanding rather than exploitation of multiple-choice structural patterns, validating that our approach teaches robust spatial reasoning capabilities that generalize across different response formats.
Multi-Backbone Generalization Validation. To demonstrate the broader applicability of our training methodology, we evaluated MVSM training procedures across multiple vision-language model architectures using the same multi-perspective spatial dataset described in Section 4 and the training configurations specified in Section 5.1. As shown in the table 6, results across three representative backbones show consistent substantial improvements: Qwen2.5-VL(3B) improved from 35.85% to 82.09% (+46.24%), InternVL-2B improved from 34.98% to 76.45% (+41.47%), and Qwen2.5-VL(7B) improved from 36.85% to 83.01% (+46.16%). The consistent performance gains across different model families and parameter scales demonstrate the architecture-agnostic effectiveness of our perspective-aware training approach. These results establish that the benefits of multi-perspective spatial training extend beyond specific model implementations, indicating robust transferability of our methodology across diverse vision-language architectures.
Figure 7: ViewSpatial-Bench Examples (Part1). Performance comparison of three models (Qwen2.5-VL(3B), GPT-4o, and MVSM) on five spatial reasoning tasks from camera perspective ( ) and person perspective ( ).